a candy distributor needs to mix a 20% fat-content chocolate with 60% fat-content chocolate to create a 200 kilograms of a 26% fat-content chocolate. How many kilograms of each kind of chocolate must they use?
Answer by richwmiller(17219)

a+b=200,
0.2*a+0.6*b=0.26*200
a=200-b
0.2*(200-b)+0.6*b=52
40-0.2b+0.6*b=52
0.4*b=12
b=30
a=200-b
a=170 kg at 20%
b=30 kg at 60%
check
0.2*170+0.6*30=0.26*200
34+18=52
52=52
ok